How would I find the domain of y = root(x-3) - root(x+3) ?
And how would I write that?
What about the domain of y = (x^2 + 8x + 12)/4throot(x+5) ?
Or y = log(2x-12) ?

Answer by stanbon(75887)

How would I find the domain of y = root(x-3) - root(x+3) ?
And how would I write that?
---
Since the radicands must not be negative, x-3 must be >=0
Domain: All Real Numbers,x, where x >= 3.
--------------------------------------------
What about the domain of y = (x^2 + 8x + 12)/4throot(x+5) ?
x+5 mult be > 0
Domain: All Real Numbers, x, where x > -5.
----------------------------------------------
Or y = log(2x-12) ?
2x-12 must be > 0
Domain: All Real Numbers, x, where x > 6.
================================
Cheeers,
Stan H.